OpenStudy (anonymous):

Complete the table for the equation x + 2y = 7, and graph the equation.

OpenStudy (anonymous):

Can someone help me out?

OpenStudy (anonymous):

Do know how to set this up

OpenStudy (anonymous):

write the equation in standart form: x+2y-x=7-x, so 2y=7-x. Divide both sides by 2. y=(7-x)/2=7/2+x/2=(1/2)*x+3.5. Now substitute any value for x and solve for y. The graph is 

OpenStudy (anonymous):

How did you come up with these points?

OpenStudy (anonymous):

First graph y=(1/2)*x Then shift this graph 3.5 units up

OpenStudy (anonymous):

Do want the points it already have in the table on my problem

OpenStudy (anonymous):

Besides you can plot points you solved and result the same graph

OpenStudy (anonymous):

You mean you will give me the table to fill in?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok. just a sec

OpenStudy (anonymous):

alright

OpenStudy (anonymous):